OpenStudy (anonymous):

Solve x = r2 - y2 for r.

mathslover (mathslover):

ok .. so first add y2 on both sides i meant : \[\large{x+y^2=r^2+y^2-y^2}\]

OpenStudy (anonymous):

Thanks

mathslover (mathslover):

my pleasure and also : \[\Huge{\mathbb{Welcome}\textbf{To}\mathbb{OpenStudy}}\]

mathslover (mathslover):

i think you did wrong some where: \[\large{x^2+y^2=r^2-y^2+y^2}\] \[\large{x^2+y^2=r^2-\cancel{y^2}+\cancel{y^2}}\] \[\large{x^2+y^2=r^2}\] \[\large{r=\sqrt{x^2+y^2}}\] your answer